Skip to main content

AAVSOnet Status, 2018-09-12

1 post / 0 new
HQA
HQA's picture
AAVSOnet Status, 2018-09-12

This has not been the best of times for the network!  Here is a list of the current status for the various telescope systems.

BSM_Berry  Fully operational.

BSM_Hamren  Camera sent in to SBIG for service.  Their prognosis is complete failure, so we will have to replace this camera.  As there are several decisions involved, BSM_Hamren will be down for the near future until we have a clear path forward.

BSM_NH  Fully operational.  We had a recent period where the pointing degraded, but have since rebuilt the pointing model.  We have new equipment that is being tested and will be installed over the next week or so.

BSM_NM  Camera sent in to SBIG for service, repaired and is being returned.  The monsoon has basically ended, so we hope to be up and running shortly.

BSM_S  Some mount problems have been resolved, so the system is now operational.

SRO  Monsoon has ended.  John Gross is going to the observatory this weekend to turn on equipment, so operations should begin shortly.

OC61  A number of problems occurred over the past month.  The mirrors were realuminized in mid-August, but new flats were not available until recently.  Some automatic processing occurred with incorrect flats.  You may get multiple "you've got data" emails as we correct this problem.  More recently, the dome motor and slit motor have failed.  The dome motor has been repaired, but the slit motor has not been replaced, and the system is down.  We hope to be back up shortly.

TMO61  Fully operational, but I haven't seen new data recently.  I'll check with the system manager and see if there is a new problem.  Monsoon is over, so operations should be resuming shortly.

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484